Please help me put this in standard form x^2 - 4x + y^2 + 8y - 16 = 0
Answer by stanbon(48580)

put this in standard form x^2 - 4x + y^2 + 8y - 16 = 0
------------
Complete the square on the x-terms and the y-terms separately:
---
x^2-4x + 4 + y^2 + 8y + 16 = 16 + 4 + 16
(x - 2)^2 + (y + 4)^2 = 36
---
[(x-2)^2)/36] + [(y+4)^2)/36] = 1
=======================================
Cheers,
Stan H.